OpenStudy (anonymous):

What is the resultant momentum of a system of two particles with these masses and velocities: 3.75kg m/s moving North and 4.2 kg moving Northwest at 2.3 m/s?

OpenStudy (anonymous):

The 3.75 body is moving at 5.5 m/s north

OpenStudy (anonymous):

 The momentum of a particle is given by the product of its mass and velocity:\[p=m\times v\] To add momentums, they have to be going in the same direction. Your best bet is to break the momentum of the second particle into its vertical and horizontal components:\[p_V=p \times sin \theta\]\[p_H=p \times cos\theta\]